OpenStudy (anonymous):

Algebra Help- I have to make a equation based upon a point and its slope. I will draw it out.

OpenStudy (anonymous):



OpenStudy (anonymous):

the equation I found for it is \[y=\frac{ 1 }{\sqrt{3} }-1\sqrt{3}x\] How do I simplify this?

OpenStudy (anonymous):

Which of the following points are on this slope? A(-2,sqrt(3)), B(0,1), C(-1/3,2/sqrt(3)) D(-sqrt(3),2sqrt(3))

OpenStudy (anonymous):

@ajprincess do you know how to solve this?

Parth (parthkohli):

Point and its slope? Point-slope form!

OpenStudy (anonymous):

@ParthKohli that part I did. But it gets ms confused algebretically solving with squareroots. How do I find all those points if they exists on the equation?

Parth (parthkohli):

What is the question exactly?

OpenStudy (anonymous):

after finding the equation, find if those points listed above are on that line or not.

Parth (parthkohli):

Plug and play.

OpenStudy (anonymous):

@parthkohli I tried, but it gets messed up. Can you help me step by step?

OpenStudy (ajprincess):

Let us substitute the first value and see. (-2,sqrt(3)), \(y=1/\sqrt3-\sqrt3x\) \(=1/\sqrt3-\sqrt3(-2)\) \( =1/\sqrt3+2\sqrt3\) \(=(1+2\sqrt3(\sqrt3)/\sqrt3\) \(=(1+2*3)/sqrt3\) \(=(1+6)/\sqrt3\) \(=7/\sqrt3\) \(7/\sqrt3\ne\sqrt3\) So this is nt a point. Similarly can u chck the others too. @Comgroupmail?

OpenStudy (anonymous):

Plug in C: \[y = 1/\sqrt{3} - \sqrt{3}(-1/3) = 1/\sqrt{3} + 1/\sqrt{3} = 2/\sqrt{3}\]C is the answer. Do it this way and you will find all the others wrong

OpenStudy (anonymous):